At one store, 5 pairs of jeans and 2 sweatshirts costs $213, while 3 pairs of jeans and 4 sweatshirts costs $195. Find the cost of one sweatshirt.
A. $21
B. $35
C. $33
D. $24

Answer by math_helper(2461)

J = price of one pair of jeans
S = price of a sweatshirt

(1) 5J+2S = 213
(2) 3J +4S = 195
(1') 10J+4S = 426 (multiplied eq(1) by 2)
-(2) 3J +4S = 195
——————————
7J = 231
J = 231/7 = 33
Plug this into (1) and solve for S (one way to do it):
S = (213 - 5J)/2 = (213 - 5*33)/2 = 24

Ans: One sweatshirt costs $24

Check:
5*33+2*24 = 165+48 = 213
3*33+4*24 = 99+96 = 195